# Learning Versatile Optimizers on a Compute Diet

### v1

@@ 1 Introduction
Deep learning has advanced remarkably over the past decade by significant improvements in architectures and training techniques, allowing us to train large neural networks on big datasets (Krizhevsky et al., 2012; Achiam et al., 2023; Brooks et al., 2024; Kirillov et al., 2023; Jumper et al., 2021). Much of this progress can be attributed to advancements in optimizers (Sutskever et al., 2013; Kingma & Ba, 2015; Loshchilov, 2017; Gupta et al., 2018; Rajbhandari et al., 2020) updating the neural networks (optimizee) over the course of training. In deep learning, the most popular optimizers are adaptive variants of stochastic gradient descent such as Adam (Kingma & Ba, 2015; Loshchilov, 2017), AdaGrad (Duchi et al., 2011), Adafactor (Shazeer & Stern, 2018), etc. These variants, although supported by some theoretical guarantees, are often heuristically developed and tested on large-scale neural network tasks (Goodfellow et al., 2016; Schmidt et al., 2021; Bottou et al., 2018; Reddi et al., 2019). In addition to being hand-designed (thus being potentially sub-optimal), these optimizers have hyperparameters that need to be tuned for every task, which is quite compute intensive (Schmidt et al., 2021; Sivaprasad et al., 2020). Additionally, recent works also use schedules (Goyal, 2017; You et al., 2017; Smith, 2017; Loshchilov & Hutter, 2016) which adjust optimizer hyperparameters dynamically throughout training. These schedules introduce additional tunable hyperparameters, making the whole training process even more computationally expensive (Metz et al., 2020b; Dahl et al., 2023).
@@
Another challenge in learned optimization is quantitative evaluation which can be used as a reliable measure to track progress in this field. Most prior work in learned optimizers (Schneider et al., 2019; Schmidt et al., 2021; Metz et al., 2022b; a) focus on loss curves or final performance of the trained models on selected tasks as the primary mode of evaluation. However, it remains unclear how to obtain a reliable aggregated performance of an optimizer across a diverse, large set of tasks, each with potentially different evaluation metrics and loss scales. Furthermore, learned optimizers often also exhibit high variance and occasional instability across tasks (Andrychowicz et al., 2016; Metz et al., 2022a; b) which can also significantly bias the evaluation results due to outliers. In this work, we seek inspiration from Reinforcement Learning (Agarwal et al., 2021), which also faces similar evaluation challenges and proposes reliable evaluation metrics which we employ to quantitatively evaluate learned optimizers. Our contributions in this work are the following:
@@ 2 Background
Unlike hand-designed optimizers, learned optimizers’ update function is parametrized by $\bm{\phi}$ learned through a meta-training process. A standard approach to meta-training involves solving a bi-level optimization problem that involves an inner problem which optimizes network parameters $\bm{\theta}$ using the learned optimizer update $\bm{\phi}$ on a sampled task and an outer problem which optimizes the learned optimizer parameters $\bm{\phi}$ based on the the feedback from the inner loop (Andrychowicz et al., 2016; Wichrowska et al., 2017; Metz et al., 2019a; 2020a; Vicol et al., 2021). Formally, given a set of optimization tasks $\mathcal{T}$, the learned optimizer parameters $\bm{\phi}$ are obtained by sampling an optimization task which consists of data distribution $\mathcal{D}$, initial network parameters $\bm{\theta}_{0}$, and a training objective $\mathcal{L}$ and solving the bi-level problem below:
@@
Here $T$ denotes the unroll length in the inner loop and $X$ denotes sampled data from $\mathcal{D}$. The outer training objective or meta-objective is based on the mean loss as formalized above or, less common, the final loss of the inner loop (Metz et al., 2019a; 2020a; 2022b). We compute meta-gradients for $\bm{\phi}$ in eq. 3 using Persistent Evolution Strategies (PES) (Vicol et al., 2021) used in prior work (Metz et al., 2022a; Harrison et al., 2022; Gärtner et al., 2023) since directly back-propagating through the inner loop in eq. 3 can lead to noisy gradient estimates especially when the unroll length $T$ is large (Metz et al., 2019a; Vicol et al., 2021). Moreover, PES also gives unbiased gradient estimates in the truncated unroll setting in which the learned optimizer is updated in truncations every few steps in the inner loop instead of updating only once after the full unroll of the inner loop. This allows frequent updates to the optimizer during meta-training thus making it less computationally expensive than full unrolls (Vicol et al., 2021; Metz et al., 2022b).
@@ 3 Celo: Compute-efficient learned optimizer
Building on the success of data augmentation strategies in deep learning (Krizhevsky et al., 2012; Cubuk et al., 2020; Chen et al., 2020; Laskin et al., 2020), we apply task augmentation during meta-training to enhance the robustness of learned optimizers (Metz et al., 2022b). This approach emulates learning dynamics across a potentially large number of tasks using a limited set of meta-training tasks (Metz et al., 2022b). Task augmentation is done by re-scaling parameters (re-parametrizing) in the inner loop. Recall that each meta-training iteration consists of applying an inner training loop on a randomly initialized, optimizee network with parameters, $\bm{\theta}_{0}$. A random scalar parameter $\tau$ is sampled in each meta-training iteration by which the optimizee network parameters at each inner training step $t$ are scaled by $\tau$ before the forward pass as $\tau\bm{\theta}_{t}$ (re-parametrization) which yields inner task loss $\mathcal{L}$. At initialization, the initial weights $\bm{\theta}_{0}$ are also scaled by a factor of $1/\tau$ in order to keep the underlying optimizee function output same. Modified Eq. 3 after task augmentation is given below with task augmentation specific changes highlighted in red:
@@
Several designs for learned optimizers have been proposed in prior work which extend this notion of schedulers beyond a scalar learning rate (Andrychowicz et al., 2016; Wichrowska et al., 2017; Almeida et al., 2021; Peebles et al., 2022; Metz et al., 2022b), see Fig. 8 in Appendix. However, in this work, we streamline the learned optimizer design and show that a learned recurrent scheduler (LSTM in our case) which outputs a scalar schedule param coupled with learned per-parameter MLP update rule generalizes well to a large-number of unseen tasks, when even meta-trained on a limited number of small-scale tasks. Our learned LSTM-based scheduler takes global loss and progress features (Metz et al., 2022b; a) as input and outputs a positive scalar $\eta_{t}$ which is obtained by linear projecting the LSTM output and raising it to the exponential. We show that this exponential form in scheduler is crucial for performance in §5. The final parameter updates are obtained by scaling the per-parameter MLP (Metz et al., 2022a; b) outputs with this scalar $\eta_{t}$ predicted by the scheduler. The pseudocode for forward pass of our optimizer is expressed concretely in Algorithm 1.
@@
We propose to view an optimization problem in two parts: how to control step-size and what rule to use for parameter updates. Current state-of-the-art hand-designed optimizers such as Adam (Kingma & Ba, 2015) and Shampoo (Gupta et al., 2018) propose a hand-designed update rule, leaving the task of controlling or modulating global learning rate and other hyperparameters up to the practitioners. Following this view, we propose to meta-train learned optimizers in a two-stage process: (1) Learn a strong param update rule (2) Freeze the learned update rule and learn a step-size scheduler. Our experiments show that two-stage training is crucial for meta-generalization and avoids meta-overfitting when training on a compute budget with fixed amount of meta-training tasks, hence outperforming single-stage training. Moreover, our simple two-stage design allows us to contrast and compare with state-of-the-art hand-designed optimizers. Our experiments show that our learned update rule outperforms hand-designed update rules like Adam and when combined with a learned scheduler, the meta-generalization performance further improves.
@@ 4 Reliably scoring optimizers at scale
Prior work in learned optimizers often resorts to showing training or validation curves as the primary means of evaluation (Schneider et al., 2019; Schmidt et al., 2021; Metz et al., 2022b; a). This is a common practice, because evaluating the performance of learned optimizers quantitatively is challenging. For example, different evaluation tasks can have different ranges and semantics of loss values. Moreover, learned optimizers also often show unstable behavior and evaluating them multiple times to reduce the variance can be infeasible, since each evaluation run involves training a neural network. Therefore, metrics such as mean or median (of loss) can be noisy.
Alternatively, an aggregate speedup can be reported, as done in prior work (Dahl et al., 2023). However, this approach requires carefully designing target scores for each task and meaningfully aggregating speedups, which becomes cumbersome for a large and diverse set of potentially changing tasks. Moreover, if the targets are set too high, the optimizers that fail to meet these targets receive a zero score, disregarding improvements between the optimizers. Conversely, setting targets too low introduces a large bias from outliers, compromising the evaluation’s reliability.
@@
Similar to the setup described above, we first compute normalized speedup $M\times N$ normalized score matrix $\bm{S}$ which is then used to compute the aggregate speedup IQM metric. We compute normalized scores as follows: For a given task, we first pick best Adam baseline out of $K$ trials which achieves lowest loss at the end of training, which consists of say, $T$ steps. Next, we check the number of steps for the optimizer being evaluated to achieve the same loss, say, $T^{\text{opt}}$ and compute normalized speedup score $s_{m,n}$ as below:
@@ 7 Related work
Learned optimizer architectures. Learned optimizer architectures can be broadly classified into three categories: (1) Flat learned optimizers (2) Hierarchical learned optimizers (3) Hybrid optimizers. Flat optimizers (Andrychowicz et al., 2016; Metz et al., 2022a; Harrison et al., 2022) use the same update rule for every parameter in the optimizee neural network without considering any additional context such as structure of the underlying optimizee network. Hence, their parameter update computation is non-hierarchical or in other words “flat”. These flat learned optimizers can be thought of as a drop-in replacement for hand-crafted update rules such as SGD or Adam. Hierarchical optimizers (Wichrowska et al., 2017; Metz et al., 2020a; 2022b; Peebles et al., 2022; Moudgil et al., 2023), unlike flat optimizers, have a hierarchical structure in their update function. They take additional meta-data of the optimizee neural network such as tensor shapes, neural graph, etc as input and utilize it in their parameter update function. These hierarchical optimizers can not only capture inter-parameter dependencies for faster optimization but also be scaled in memory-efficient manner without significantly increasing the runtime overhead since their runtime cost scale sub-linearly with the number of parameters due to hierarchy Metz et al. (2022b; a). Hybrid optimizers (Almeida et al., 2021; Knyazev et al., 2024; Kristiansen et al., 2024) combine hand-crafted and learned modules in their update function. Using statistics of the underlying optimization task such as loss, gradient norms, momentum, etc, learned modules in hybrid optimizers control hyperparameters of the hand-crafted update rules such as SGD (Kristiansen et al., 2024), Adam (Metz et al., 2020b; Kristiansen et al., 2024) or LAMB (Almeida et al., 2021). In this work, we focus on enhancing the meta-generalization capabilities of fully learned optimizers while maintaining memory and compute efficiency for practical use. Hence, our proposed approach, Celo, has a hierarchical architecture.
@@
Meta-generalization of learned optimizers. Several learned optimizers have been proposed for specific settings such as physics-based motion reconstruction (Gärtner et al., 2023), reinforcement learning (Goldie et al., 2024), material design (Merchant et al., 2021), etc. Although these approaches are quite performant, their applicability to broad range of Machine Learning (ML) tasks that practitioners care about is limited. Since meta-training a learned optimizer for each setting is computationally expensive, optimizers that meta-generalize to multiple tasks after being meta-trained are more desirable. A few recent works have explored this direction.  Almeida et al. (2021) proposed a hybrid optimizer with LAMB update rule that is trained on a small-scale tasks and generalize to 3 orders of magnitude higher tasks. VeLO (short for “Versatile Learned Optimizer”) is a fully learned optimizer, meta-trained using an extensive computational budget of 4000 TPU-months across diverse ML tasks. VeLO has demonstrated the ability to generalize effectively to new, unseen tasks and successfully optimize neural networks with parameters up to 600M. In this work, we propose a recipe to meta-train fully learned optimizers on a limited compute budget that meta-generalize to wide range of unseen tasks.
@@
Benchmarking optimizers. A few works have attempted to benchmark and compare optimizers (Schneider et al., 2019; Schmidt et al., 2021; Dahl et al., 2023) on standard ML  Schneider et al. (2019) proposes DeepOBS benchmark that contains a set of realistic 8 ML tasks and compares standard (hand-crafted) optimizers using training curves, test curves and learning rate sensitivity analysis. Schmidt et al. (2021) benchmarks 15 hand-crafted optimizers on 8 ML tasks from DeepOBS by varying tuning budget and learning rate schedules, leading to more than 50K optmization runs from all the combinations. The optimizers are compared relative to each other by using final loss/test accuracy.  Dahl et al. (2023) proposes another benchmark that evaluates standard optimizers on 8 large-scale ML-Perf tasks and a methodology to score optimizers by using speedup. Speedups for each task are computed using fixed task-specific metric targets that are hand-designed. In this work, we propose evaluation metrics that can capture aggregate performance of an optimizer using both speedup and final performance criteria. Unlike prior work, our evaluation metrics are specifically designed to be robust to outliers especially when the number of evaluation tasks is large and doesn’t require hand-designing any cut-offs/thresholds. Hence, our proposed evaluation methodology is scalable with respect to number of evaluation tasks.
8 Scope and limitations

### v2

# Celo: Training Versatile Learned Optimizers on a Compute Diet

@@ 1 Introduction
Deep learning has advanced remarkably over the past decade by significant improvements in architectures and training techniques, allowing us to train large neural networks on big datasets (Krizhevsky et al., 2012; Achiam et al., 2023; Brooks et al., 2024; Kirillov et al., 2023; Jumper et al., 2021). Much of this progress can be attributed to advancements in optimizers (Sutskever et al., 2013; Kingma & Ba, 2015; Loshchilov, 2017; Gupta et al., 2018; Rajbhandari et al., 2020) updating the neural networks (optimizee) over the course of training. In deep learning, the most popular optimizers are adaptive variants of stochastic gradient descent such as Adam (Kingma & Ba, 2015; Loshchilov, 2017), AdaGrad (Duchi et al., 2011), Adafactor (Shazeer & Stern, 2018), etc. These variants, although supported by some theoretical guarantees, are often heuristically developed and tested on large-scale neural network tasks (Goodfellow et al., 2016; Schmidt et al., 2021; Bottou et al., 2018; Reddi et al., 2019). In addition to being hand-designed (thus being potentially suboptimal), these optimizers have hyperparameters that need to be tuned for every task, which is quite compute intensive (Schmidt et al., 2021; Sivaprasad et al., 2020). Additionally, recent works also use schedules (Goyal, 2017; You et al., 2017; Smith, 2017; Loshchilov & Hutter, 2016) which adjust optimizer hyperparameters dynamically throughout training. These schedules introduce additional tunable hyperparameters, making the whole training process even more computationally expensive (Metz et al., 2020b; Dahl et al., 2023).
@@
Another challenge in learned optimization is quantitative evaluation, which can be used as a reliable measure to track progress in this field. Most prior work in learned optimizers (Schneider et al., 2019; Schmidt et al., 2021; Metz et al., 2022b; a) focus on loss curves or final performance of the trained models on selected tasks as the primary mode of evaluation. However, it remains unclear how to obtain a reliable aggregated performance of an optimizer across a diverse, large set of tasks, each with potentially different evaluation metrics and loss scales. Furthermore, learned optimizers often also exhibit high variance and occasional instability across tasks (Andrychowicz et al., 2016; Metz et al., 2022a; b) which can also significantly bias the evaluation results due to outliers. In this work, we seek inspiration from Reinforcement Learning (Agarwal et al., 2021), which also faces similar evaluation challenges and proposes reliable evaluation metrics which we employ to quantitatively evaluate learned optimizers. Our contributions in this work are the following:
@@ 2 Background
Unlike hand-designed optimizers, learned optimizers’ update function is parametrized by $\bm{\phi}$ learned through a meta-training process. A standard approach to meta-training involves solving a bi-level optimization problem that involves an inner problem which optimizes network parameters $\bm{\theta}$ using the learned optimizer update $\bm{\phi}$ on a sampled task and an outer problem which optimizes the learned optimizer parameters $\bm{\phi}$ based on the feedback from the inner loop (Andrychowicz et al., 2016; Wichrowska et al., 2017; Metz et al., 2019a; 2020a; Vicol et al., 2021). Formally, given a set of optimization tasks $\mathcal{T}$, the learned optimizer parameters $\bm{\phi}$ are obtained by sampling an optimization task which consists of data distribution $\mathcal{D}$, initial network parameters $\bm{\theta}_{0}$, and a training objective $\mathcal{L}$ and solving the bi-level problem below:
@@
Here $T$ denotes the unroll length in the inner loop and $X$ denotes sampled data from $\mathcal{D}$. The outer training objective or meta-objective is based on the mean loss as formalized above or, less common, the final loss of the inner loop (Metz et al., 2019a; 2020a; 2022b). We compute meta-gradients for $\bm{\phi}$ in eq. 3 using Persistent Evolution Strategies (PES) (Vicol et al., 2021) used in prior work (Metz et al., 2022a; Harrison et al., 2022; Gärtner et al., 2023) since directly back-propagating through the inner loop in eq. 3 can lead to noisy gradient estimates, especially when the unroll length $T$ is large (Metz et al., 2019a; Vicol et al., 2021). Moreover, PES also gives unbiased gradient estimates in the truncated unroll setting in which the learned optimizer is updated in truncations every few steps in the inner loop instead of updating only once after the full unroll of the inner loop. This allows frequent updates to the optimizer during meta-training, thus making it less computationally expensive than full unrolls (Vicol et al., 2021; Metz et al., 2022b).
@@ 3 Celo: Compute-efficient learned optimizer
Building on the success of data augmentation strategies in deep learning (Krizhevsky et al., 2012; Cubuk et al., 2020; Chen et al., 2020b; Laskin et al., 2020), we apply task augmentation during meta-training to enhance the robustness of learned optimizers (Metz et al., 2022b). This approach emulates learning dynamics across a potentially large number of tasks using a limited set of meta-training tasks (Metz et al., 2022b). Task augmentation is done by re-scaling parameters (re-parametrizing) in the inner loop. Recall that each meta-training iteration consists of applying an inner training loop on a randomly initialized, optimizee network with parameters, $\bm{\theta}_{0}$. A random scalar parameter $\tau$ is sampled in each meta-training iteration by which the optimizee network parameters at each inner training step $t$ are scaled by $\tau$ before the forward pass as $\tau\bm{\theta}_{t}$ (re-parametrization) which yields inner task loss $\mathcal{L}$. At initialization, the initial weights $\bm{\theta}_{0}$ are also scaled by a factor of $1/\tau$ in order to keep the underlying optimizee function output same. Modified Eq. 3 after task augmentation is given below with task augmentation specific changes highlighted in red:
@@
Several designs for learned optimizers have been proposed in prior work which extend this notion of schedulers beyond a scalar learning rate (Andrychowicz et al., 2016; Wichrowska et al., 2017; Almeida et al., 2021; Peebles et al., 2022; Metz et al., 2022b), see Fig. 8 in Appendix. However, in this work, we streamline the learned optimizer design and show that a learned recurrent scheduler (LSTM in our case) which outputs a scalar schedule parameter coupled with learned per-parameter MLP update rule generalizes well to a large-number of unseen tasks, when even meta-trained on a limited number of small-scale tasks. Our learned LSTM-based scheduler takes global loss and progress features (Metz et al., 2022b; a) as input and outputs a positive scalar $\eta_{t}$ which is obtained by linearly projecting the LSTM output and raising it to the exponential. We show that this exponential form in scheduler is crucial for performance in §5. The final parameter updates are obtained by scaling the per-parameter MLP (Metz et al., 2022a; b) outputs with this scalar $\eta_{t}$ predicted by the scheduler. The pseudocode for forward pass of our optimizer is expressed concretely in Algorithm 1.
@@
We propose to view an optimization problem in two parts: how to control step-size and what rule to use for parameter updates. Current state-of-the-art hand-designed optimizers such as Adam (Kingma & Ba, 2015) and Shampoo (Gupta et al., 2018) propose a hand-designed update rule, leaving the task of controlling or modulating global learning rate and other hyperparameters up to the practitioners. Following this view, we propose to meta-train learned optimizers in a two-stage process: (1) Learn a strong parameter update rule (2) Freeze the learned update rule and learn a step-size scheduler. Our experiments show that two-stage training is crucial for meta-generalization and avoids meta-overfitting when training on a compute budget with fixed amount of meta-training tasks, hence outperforming single-stage training. Moreover, our simple two-stage design allows us to contrast and compare with state-of-the-art hand-designed optimizers. Our experiments show that our learned update rule outperforms hand-designed update rules like Adam and when combined with a learned scheduler, the meta-generalization performance further improves.
@@ 4 Reliably scoring optimizers at scale
Prior work in learned optimizers often resorts to showing training or validation curves as the primary means of evaluation (Schneider et al., 2019; Schmidt et al., 2021; Metz et al., 2022b; a). This is a common practice, because evaluating the performance of learned optimizers quantitatively is challenging. For example, different evaluation tasks can have different ranges and semantics of loss values. Moreover, learned optimizers also often show unstable behaviour and evaluating them multiple times to reduce the variance can be infeasible, since each evaluation run involves training a neural network. Therefore, metrics such as mean or median (of loss) can be noisy. Alternatively, an aggregate speedup can be reported, as done in prior work (Dahl et al., 2023). However, this approach requires carefully designing target scores for each task and meaningfully aggregating speedups, which becomes cumbersome for a large and diverse set of potentially changing tasks. Moreover, if the targets are set too high, the optimizers that fail to meet these targets receive a zero score, disregarding improvements between the optimizers. Conversely, setting targets too low introduces a large bias from outliers, compromising the evaluation’s reliability.
@@
Similar to the setup described above, we first compute a normalized speedup $M\times N$ normalized score matrix $\bm{S}$ which is then used to compute the aggregate speedup IQM metric. We compute normalized scores as follows: For a given task, we first pick the best Adam baseline out of $K$ trials which achieves the lowest loss at the end of training, which consists of say, $T$ steps. Next, we check the number of steps for the optimizer being evaluated to achieve the same loss, say, $T^{\text{opt}}$ and compute normalized speedup score $s_{m,n}$ as below:
@@ 7 Related work
Learned optimizer architectures. Learned optimizer architectures can be broadly classified into three categories: (1) Flat learned optimizers (2) Hierarchical learned optimizers (3) Hybrid optimizers. Flat optimizers (Andrychowicz et al., 2016; Chen et al., 2020a; Metz et al., 2022a; Harrison et al., 2022) use the same update rule for every parameter in the optimizee neural network without considering any additional context such as structure of the underlying optimizee network. Hence, their parameter update computation is non-hierarchical or in other words “flat”. These flat learned optimizers can be thought of as a drop-in replacement for hand-crafted update rules such as SGD or Adam. Hierarchical optimizers (Wichrowska et al., 2017; Metz et al., 2020a; 2022b; Peebles et al., 2022; Moudgil et al., 2023), unlike flat optimizers, have a hierarchical structure in their update function. They take additional meta-data of the optimizee neural network such as tensor shapes, neural graph (Kofinas et al., 2024), etc as input and utilize it in their parameter update function. These hierarchical optimizers can not only capture inter-parameter dependencies for faster optimization but also be scaled in memory-efficient manner without significantly increasing the runtime overhead since their runtime cost scale sub-linearly with the number of parameters due to hierarchy (Metz et al., 2022b; a). Hybrid optimizers (Almeida et al., 2021; Jang et al., 2023; Knyazev et al., 2024; Kristiansen et al., 2024) combine hand-crafted and learned modules in their update function. Using statistics of the underlying optimization task such as loss, gradient norms, momentum, etc, learned modules in hybrid optimizers control hyperparameters of the hand-crafted update rules such as SGD (Kristiansen et al., 2024), Adam (Metz et al., 2020b; Kristiansen et al., 2024) or LAMB (Almeida et al., 2021). Jang et al. (2023); Knyazev et al. (2024) alternate between Adam updates and updates performed by a learned function taking only the past parameters as input showing promising results, but limited scalability. In this work, we focus on enhancing the meta-generalization capabilities of fully learned optimizers while maintaining memory and compute efficiency for practical use. Hence, our proposed approach, Celo, has a hierarchical architecture.
@@
Meta-generalization of learned optimizers. Several learned optimizers have been proposed for specific settings such as physics-based motion reconstruction (Gärtner et al., 2023), reinforcement learning (Goldie et al., 2024), material design (Merchant et al., 2021), etc. Although these approaches are quite performant, their applicability to broad range of Machine Learning (ML) tasks that practitioners care about is limited. Since meta-training a learned optimizer for each setting is computationally expensive, optimizers that meta-generalize to multiple tasks after being meta-trained are more desirable. A few recent works have explored this direction.  Almeida et al. (2021) proposed a hybrid optimizer with LAMB update rule that is trained on small-scale tasks and generalize to 3 orders of magnitude higher tasks. VeLO (short for “Versatile Learned Optimizer”) is a fully learned optimizer, meta-trained using an extensive computational budget of 4000 TPU-months across diverse ML tasks. VeLO has demonstrated the ability to generalize effectively to new, unseen tasks and successfully optimize neural networks with parameters up to 600M. In this work, we propose a recipe to meta-train fully learned optimizers on a limited compute budget that meta-generalize to wide range of unseen tasks.  Thérien et al. (2024) recently proposed another approach to tackle meta-generalization by developing learned optimizers within the Maximal Update Parametrization (Yang et al., 2021) framework. This enables meta-generalization to wider and deeper models as well as to longer unrolls of a meta-training task. Their approach is complementary to ours, so combining it with our recipe is a promising direction to further improve meta-generalization.
@@
Benchmarking optimizers. A few works have attempted to benchmark and compare optimizers (Schneider et al., 2019; Schmidt et al., 2021; Dahl et al., 2023) on standard ML. Schneider et al. (2019) proposed the DeepOBS benchmark that contains a set of realistic 8 ML tasks and compares standard (hand-crafted) optimizers using training curves, test curves and learning rate sensitivity analysis. Schmidt et al. (2021) benchmarked 15 hand-crafted optimizers on 8 ML tasks from DeepOBS by varying tuning budget and learning rate schedules, leading to more than 50K optimization runs from all the combinations. The optimizers are compared relative to each other by using final loss/test accuracy.  Dahl et al. (2023) proposed another benchmark that evaluates standard optimizers on 8 large-scale ML-Perf tasks and a methodology to score optimizers by using speedup. Speedups for each task are computed using fixed task-specific metric targets that are hand-designed. In this work, we propose evaluation metrics that can capture aggregate performance of an optimizer using both speedup and final performance criteria. Unlike prior work, our evaluation metrics are specifically designed to be robust to outliers, especially when the number of evaluation tasks is large and doesn’t require hand-designing any cut-offs/thresholds. Hence, our proposed evaluation methodology is scalable with respect to number of evaluation tasks.
8 Limitations and future work